OpenStudy (90890908967777878888):

It takes Nadia 4/9 h to complete her Language Arts homework and 2/9 h to complete her mathematics homework. Which takes longer? How much longer?

OpenStudy (anonymous):

What is longest \[4/9h\] or \[2/9h\]?

OpenStudy (anonymous):

Language Arts, 2/9 hour longer

OpenStudy (90890908967777878888):

4/9

OpenStudy (anonymous):

Since they have the same denominator (bottom number), it's simple to tell which is larger based on the numerator (top number) To get the time difference, just subtract: 4/9 - 2/9 = 2/9

OpenStudy (90890908967777878888):

so how i explain

OpenStudy (90890908967777878888):

@porkbuns how i explain

OpenStudy (90890908967777878888):

it 2/9

OpenStudy (anonymous):

Since 4/9 and 2/9 have the same denominator (bottom #), you can tell which is larger by the numerator (top #), 4>2 (Language Arts takes longer) To get the time difference, just subtract 2/9 from 4/9: 4/9 - 2/9 = 2/9 hour longer

OpenStudy (90890908967777878888):

so how i explain